Jefferson Lab HALL A Logbook for August 10 - All EXCEPT Automatic Entries

  • Main INDEX

    Tuesday

  • 331984 08/31/10 23:08 A. Camsonne ARS board 0xa0000 Firmware updated
  • Monday

  • 331961 08/30/10 18:41 fansler Warm Return
  • 331960 08/30/10 18:28 A. Camsonne ARS read-out board list
  • Sunday

  • 331954 08/29/10 16:09 A. Camsonne Split size set back to 2000 MBytes
  • 331951 08/29/10 15:32 A. Camsonne Led run 6134 block 0
  • 331949 08/29/10 15:25 A. Camsonne Long LED run 6133 ( 20 splits ! )
  • 331937 08/29/10 10:32 A. Camsonne trigger_left.map modified
  • 331936 08/29/10 10:27 atrig Trigger-Download
  • 331935 08/29/10 10:25 atrig Trigger-Download
  • 331934 08/29/10 10:24 atrig Trigger-Download
  • 331933 08/29/10 10:22 Bogdan / Ale Camac modules moved in Left HRS
  • Saturday

  • 331931 08/28/10 20:57 A. Camsonne ARS pedestal run 6127
  • 331929 08/28/10 20:41 A. Camsonne LED run scan column 0 to 9 - Run 6126
  • 331928 08/28/10 20:35 Carlos DVCS HV & LV off
  • 331906 08/28/10 17:08 A. Camsonne Run 6115 LED run
  • 331898 08/28/10 14:48 A. Camsonne LED trigger
  • Friday

  • 331897 08/27/10 17:18 siqbal Peaks in L-arm S2 R-PMT ADCs 12-15
  • 331896 08/27/10 15:18 Butler/Fansl Cold Return
  • Thursday

  • 331887 08/26/10 15:45 Butler/Fansl Cold Return
  • 331886 08/26/10 10:23 A. Camsonne HV control copied from adaqh1 to adev account
  • 331885 08/26/10 07:17 J. Butler(Te DVCS Cables
  • Wednesday

  • 331884 08/25/10 06:57 fansler Warm Return
  • Tuesday

  • 331878 08/24/10 18:22 A. Camsonne Run 6102 with DVCS crates
  • Monday

  • 331865 08/23/10 17:14 A. Camsonne Left HRS HV on for cosmics data taking
  • 331863 08/23/10 17:10 A. Camsonne Config DVCS2Crates working
  • Sunday

  • 331850 08/22/10 18:27 Vincent Abbo ARSBoard in DVCS crate
  • 331847 08/22/10 17:31 A. Camsonne Second DVCS crate added to ts_scaler.crl
  • 331845 08/22/10 16:53 A. Camsonne This was for Left Arm switch 192 subnetwork
  • 331844 08/22/10 16:44 A. Camsonne Port 2 and 5 disabled / non functionning
  • Thursday

  • 331843 08/19/10 17:34 Jack Left Stack is back in
  • 331842 08/19/10 13:47 fansler Oscillator replacement interval
  • 331841 08/19/10 11:48 Fansler/Ewin N2 Calibration
  • 331840 08/19/10 11:07 withers iochacs adc database running
  • Tuesday

  • 331839 08/17/10 21:24 A. Camsonne IOC iochacs on the network
  • 331838 08/17/10 15:41 Fansler/Butl LQ3 Transister Banks
  • Monday

  • 331836 08/16/10 19:51 ejensen Left arm GC is okay, problem resides on level of trigger electronics
  • Friday

  • 331835 08/13/10 14:06 Fansler RDI Protection circuit Newport Controllers
  • 331834 08/13/10 08:25 Fansler RQ3 Mass flow
  • Thursday

  • 331831 08/12/10 19:34 A. Camsonne Run 6078 - High rate test run
  • 331823 08/12/10 17:06 A. Camsonne Blaster test result with Gigabit ethernet
  • 331812 08/12/10 16:20 A. Camsonne Network upgrade complete
  • 331810 08/12/10 09:59 A. Camsonne Network upgrade in progress - No network in the Hall
  • Wednesday

  • 331808 08/11/10 16:00 Fansler Mass flow
  • 331807 08/11/10 11:40 Fansler RD1 lead flows
  • Tuesday

  • 331806 08/10/10 22:10 A. Camsonne Run 6069 low rate ARS
  • 331802 08/10/10 18:03 A. Camsonne Low rate ARS test run - Run 6067
  • 331796 08/10/10 16:56 A. Camsonne Long High rate ARS test run - Run 6064
  • 331782 08/10/10 15:23 A. Camsonne Run 6058 with 7 ARS board
  • 331779 08/10/10 14:52 J. Butler(te Newport Controllers
  • Monday

  • 331776 08/09/10 16:36 R. Michaels ts_scaler.crl update for helicity
  • 331765 08/09/10 16:08 J. Butler(te Newport Controllers
  • 331763 08/09/10 16:06 J. Butler(te Bogies Cameras
  • 331762 08/09/10 16:05 fansler LQ1
  • 331745 08/09/10 10:46 Fansler RQ1 lead flow
  • Sunday

  • 331741 08/08/10 18:50 A. Camsonne HRS cosmics run with DVCS ARS crate - run 6038
  • 331737 08/08/10 18:27 A. Camsonne TS crl file modified to add DVCS crate
  • 331736 08/08/10 16:21 A. Camsonne DVCS Test crate connected to Trigger Supervisor
  • Saturday

  • 331732 08/07/10 16:32 A. Camsonne DVCS test crate power supply control on hatsv40 2003
  • 331725 08/07/10 14:51 A. Camsonne ARS in test crate
  • 331722 08/07/10 11:35 A. Camsonne Moved SIS scaler from Left Arm Happex crate to DVCS test crate
  • Friday

  • 331721 08/06/10 17:36 A. Camsonne Copy of crl files from test computer in dvcs2 directory in a-onl account
  • 331720 08/06/10 17:08 A. Camsonne DVCS test crate in Left HRS
  • 331719 08/06/10 14:00 Jack Left Arm S1 and Gas Cherenkov
  • Thursday

  • 331715 08/05/10 13:08 fansler LQ3
  • 331714 08/05/10 13:07 fansler Left Dipole lead flows
  • 331713 08/05/10 11:38 S.Abrahamyan LHRS vdc check
  • 331704 08/05/10 09:41 fansler LQ2/LQ3
  • Wednesday

  • 331703 08/04/10 18:56 S.Abrahamyan VDC Left Arm
  • 331694 08/04/10 17:27 Jack (Techs) Left Q1 Danfysik power supply control board replaced
  • Tuesday

    
    

    Monday

  • 331688 08/02/10 17:10 S.Abrahamyan Left Arm VDC check